[figure: 9.7 cm]

SPECIFIC HEAT OF La_{1−x}Sr_{x}MnO_{3} COMPOUNDS

A. Szewczyk^{a}, K. Piotrowski^{a}, M.U. Gutowska^{a}, B. Da̧browski^{b}, Z. Bukowski^{b}, M. Gutowski^{a}, D. Żymierska^{a}

aInstitute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, Warsaw, Poland b Northern Illinois University, DeKalb, USA

Strong exchange interactions and spin-charge-lattice couplings lead to complex properties of La_{1−x}Sr_{x}MnO_{3} for x ≈ 0.12. On cooling, structural transition appears near 270 K, followed by coupled ferromagnetic-structural transition at 140 K. Using quasiadiabatic method, the specific heat of dense polycrystalline samples with x = 0.11 and x = 0.135 was measured over the temperature range 4–280 K. Anomalies connected with structural and magnetic phase transitions were observed and analyzed. Because a poor thermal conductivity of the material, very long sample-calorimeter relaxation time was observed and the conventional method of determining heat capacity could not be used. Instead, a curve fitting method proposed in [?] was adapted and modified by applying the nearly equal ripple filters [?] to smooth the measured time dependences of temperature.

References

[1] Jih Shang Hwang, Kai Jan Lin, Cheng Tien Rev. Sci. Instrum. 68(1) 94 (1997)

[2] J.F. Kaiser and W.A. Reed Rev. Sci. Instrum. 48(11) 1447 (1977)

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szewczyk

Address for correspondence : Institute of Physics, Polish Academy of Sciences, ON-3.5 Al. Lotników 32/46, Warsaw, Poland

Email address : szewc@ifpan.edu.pl